OpenStudy (anonymous):

Simplify the expression: (c/3) ÷ (5/6)

OpenStudy (anonymous):

(c/3)/(5/6)

OpenStudy (anonymous):

@FibonacciChick666

OpenStudy (fibonaccichick666):

so, when you are dividing by a fraction what is the rule?

OpenStudy (anonymous):

um i dont remeber right now

OpenStudy (fibonaccichick666):

systematically, what do you do? It is very important

OpenStudy (fibonaccichick666):

you....chant it! Flip the second fraction and.... MULTIPLY!

OpenStudy (anonymous):

really i didnt know that im not very good at math

OpenStudy (fibonaccichick666):

yessss It is important. Memorize the chant and never ever say you are bad at math

OpenStudy (anonymous):

so i multiply 5*6

OpenStudy (fibonaccichick666):

you have \[\frac{c}{3} \div\frac{5}{6}\]

OpenStudy (anonymous):

so \[\frac{ c }{ 3 } \div \frac{ 6 }{ 5 }\]

OpenStudy (anonymous):

and cross-multiply

OpenStudy (anonymous):

right

OpenStudy (fibonaccichick666):

no, just change the division symbol in what you typed to multiplication

OpenStudy (fibonaccichick666):

cross multiplication is used when you have an equals sign

OpenStudy (anonymous):

ok so what would it look like

OpenStudy (fibonaccichick666):

\[\frac{c}{3}\times\frac{6}{5}\]

OpenStudy (fibonaccichick666):

the second fraction is flipped and then we multiply

OpenStudy (fibonaccichick666):

just like normal

OpenStudy (anonymous):

okay so \[\frac{ 6c }{ 15 }\]

OpenStudy (fibonaccichick666):

yep, now reduce

OpenStudy (anonymous):

2c/5 @FibonacciChick666

OpenStudy (fibonaccichick666):

yeppers

OpenStudy (anonymous):

thanks i guess i understand it know

OpenStudy (fibonaccichick666):

yep, just remember, flip the second fraction and MULTIPLY! (I learned that chant in 4th grade and still use it now that I have nearly finished a math BS)